OpenStudy (anonymous):

In two years, Linda will be half as old as her mother. Her mother is 38 years old now. Let L=The age of Linda now, which formula will equal Lindas current age? 2(L+2)=38 or (L+2)=(38+2) or 2L+2=38 or 2(L+2)=(38+2)

OpenStudy (anonymous):

1st option

OpenStudy (anonymous):

Thanks

OpenStudy (lgbasallote):

in two years..linda's age will be L+2 let m = mother's age = 38 in two years mother will be (38+2)<---half of that is linda's age in the future so.. \(\large L+2 = \frac{(38+2)}{2}\) seems it's not in the option so we manipulate.. 2(L+2) = (38+2)

OpenStudy (anonymous):

Thhanks:)

OpenStudy (anonymous):

i meant first from the last row lol

OpenStudy (anonymous):

Haha yeah:)